Grade 12Algebra

Solve the following equations:
4 = 5(p – 2) …..........

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us divide both the side by 5,
= 4/5 = (5(p–2))/5
= 4/5 = p -2
By transposing - 2 from RHS to LHS it becomes 2
= (4/5) + 2 = p
= (4 + 10)/ 5 = p
= p = 14/5

Thanks